A 2.197 g sample of a compound containing only carbon, hydrogen, and oxygen is burned in an excess of dioxygen, producing 4.197 g CO2 and 2.578 g H2O. What is the empirical formula of the compound?
1
Expert's answer
2020-06-25T07:44:41-0400

First we calculate the masses of carbon and hydrogen from their oxide (CO2 and H2O).

Mass of C = 4.197 g  CO2 × 12 g C/ 44 g CO2

                = 1.14 g

Mass of H = 2.578 g  H2O × 2 g H/ 18 g H2O

                = 0.286 g

Mass of C + Mass of H = 1.426 g

This is less than the mass of sample. The missing mass must be caused by O.

Mass of O =( 2.197-1.426)g = 0.771 g

C: H: O =1.14/12 : 0.28/1 : 0.771/16

            = 0.095 : 0.286 : 0.048

            = 2 : 6 : 1

             The empirical formula of the compound is C2H6O  (Answer)